OpenStudy (swaggyswags):

The function p(x) = 2x +1 determines how many pizzas need to be purchased for an after school meeting, where x is the number of students at the meeting. The club leader uses r(p(x)) to find the amount of money to bring for the pizza purchase. The function r(x) = 4x − 3. Solve for how much money to bring when there are 9 students in the meeting. 19 33 73 84

OpenStudy (swaggyswags):

@Will.H

OpenStudy (will.h):

73 is your answer do you need explanation?

OpenStudy (swaggyswags):

Yeah, just a quick one if it's all good with you

OpenStudy (will.h):

oh explantation lol okay first you plug in x = 9 in the function p(x) and then the value would be the new value of x in the function r(x) p(x) = 2 * 9 + 1 = 19 p(rx) = 4 * 19 - 3= 73

OpenStudy (swaggyswags):

Alright sweet, thanks :)

OpenStudy (swaggyswags):

New question?

OpenStudy (will.h):

shoot

OpenStudy (swaggyswags):

Solve the following for y: −6x + 2y = 8 y = −6x + 8 y = 6x + 8 y = −3x + 4 y = 3x + 4

OpenStudy (will.h):

-6x + 2y = 8 2y = 6x + 8 y = 3x + 4

OpenStudy (swaggyswags):

Thanks dude, I appreciate the help :)

OpenStudy (will.h):

you welcome :)